OpenStudy (anonymous):

HELP ME PLZ!!!!!!! If (6y-11)(6y+11)=ay^2-b, what is the value of a?

rebeccaxhawaii (rebeccaxhawaii):

foil

OpenStudy (anonymous):

?

OpenStudy (anonymous):

6y - 11) • (6y + 11) - (y2a - b) = 0 Step 2 : Equation at the end of step 2 : -y2a + 36y2 + b - 121 = 0 Step 3 : Solving a Single Variable Equation : 3.1 Solve -y2a+36y2+b-121 = 0 In this type of equations, having more than one variable (unknown), you have to specify for which variable you want the equation solved.

OpenStudy (raffle_snaffle):